OpenStudy (anonymous):

I'm not really sure what to do on this problem. Find the solution of the differential equation y'=5x^4 e^(-x^5), such that y(0)=1

OpenStudy (anonymous):

try to set \(u=-x^5\) first

OpenStudy (anonymous):

yeah, i believe so.

OpenStudy (dumbcow):

its in the form e^-u du when u = x^5 du = 5x^4

OpenStudy (anonymous):

set \(u=x^5\) so, we get \[y=\int e^{-u}du=-e^{-u}+C=-e^{-x^5}+C\] set \(x=0\) then \[y(0)=-e^{-0^5}+C\\ 1 = -e^1+C\\ C=1+e\] i think the result is obvius by now

OpenStudy (dumbcow):

hmm \[e^{0^{5}} = e^{0} = 1\]

OpenStudy (anonymous):

oh sorry :)

OpenStudy (anonymous):

it must be C=2